OpenStudy (anonymous):

Express the quotient in simplest form 2/3 ÷ 1/6

jimthompson5910 (jim_thompson5910):

When you divide 2 fractions, you flip the second fraction and multiply \[\Large \frac{2}{3} \div \frac{1}{6} = \frac{2}{3} \times \frac{6}{1}\] do you see how to finish?

OpenStudy (anonymous):

So = 12/3

jimthompson5910 (jim_thompson5910):

then you reduce that to get 12/3 = 4